# Spatial Sign based Principal Component Analysis for High Dimensional Data

### v1

@@ 1 Introduction
PCA (Principal Component Analysis) is a widely used statistical method for data dimensionality reduction. It transforms high-dimensional data into a lower-dimensional space while preserving as much of the original data’s variability as possible. PCA achieves this by identifying the directions of maximum variance in the data, known as principal components, and projecting the data onto these directions. This process removes redundant information and noise, making the data easier to handle and visualize. PCA is commonly applied in fields such as machine learning, image processing, and finance, where dealing with high-dimensional data is common.
@@
Notation: Here we use the same notations as Han and Liu (2018). Let $\mathbf{M}=\left[\mathbf{M}_{jk}\right]\in\mathbb{R}^{d\times d}$ be a symmetric matrix and $\boldsymbol{v}=\left(v_{1},\ldots,v_{d}\right)^{T}\in\mathbb{R}^{d}$ be a vector. We denote $\boldsymbol{v}_{I}$ to be the subvector of $\boldsymbol{v}$ whose entries are indexed by a set $I$, and $\mathbf{M}_{I,J}$ to be the submatrix of $\mathbf{M}$ whose rows are indexed by $I$ and columns are indexed by $J$. We denote $\operatorname{supp}(v):=\left\{j:v_{j}\neq 0\right\}$. For $0<q<\infty$, we define the $\ell_{q}$ and $\ell_{\infty}$ vector norms as $\|\boldsymbol{v}\|_{q}:=\left(\sum_{i=1}^{d}\left|v_{i}\right|^{q}\right)^{1/q}$ and $\|\boldsymbol{v}\|_{\infty}:=\max_{1\leq i\leq d}\left|v_{i}\right|$. We denote $\|\boldsymbol{v}\|_{0}:=\operatorname{card}(\operatorname{supp}(\boldsymbol{v}))$. We define the matrix entry-wise maximum value and Frobenius norms as $\|\mathbf{M}\|_{\text{max }}:=\max\left\{\left|\mathbf{M}_{ij}\right|\right\}$ and $\|\mathbf{M}\|_{\mathrm{F}}=\left(\sum\mathbf{M}_{jk}^{2}\right)^{1/2}$. Let $\lambda_{j}(\mathbf{M})$ be the $j$ th largest eigenvalue of $\mathbf{M}$. If there are ties, $\lambda_{j}(\mathbf{M})$ is any one of the eigenvalues such that any eigenvalue larger than it has rank smaller than $j$, and any eigenvalue smaller than it has rank larger than $j$. Let $\boldsymbol{u}_{j}(\mathbf{M})$ be any unit vector $\boldsymbol{v}$ such that $\boldsymbol{v}^{T}\mathbf{M}\boldsymbol{v}=\lambda_{j}(\mathbf{M})$. Without loss of generality, we assume that the first nonzero entry of $\boldsymbol{u}_{j}(\mathbf{M})$ is positive. We denote $\|\mathbf{M}\|_{2}$ to be the spectral norm of $\mathbf{M}$ and $\mathbb{S}^{d-1}:=\left\{\boldsymbol{v}\in\mathbb{R}^{d}:\|\boldsymbol{v}\|_{2}=1\right\}$ to be the $d$-dimensional unit sphere. We define the restricted spectral norm $\|\mathbf{M}\|_{2,s}:=\sup_{\boldsymbol{v}\in\mathbb{S}^{d-1},\|\boldsymbol{v}\|_{0}\leq s}\left|\boldsymbol{v}^{T}\mathbf{M}\boldsymbol{v}\right|$, so for $s=d$, we have $\|\mathbf{M}\|_{2,s}=\|\mathbf{M}\|_{2}$. We denote $f(\mathbf{M})$ to be the matrix with entries $[f(\mathbf{M})]_{jk}=f\left(\mathbf{M}_{jk}\right)$. We denote $\operatorname{diag}(\mathbf{M})$ to be the diagonal matrix with the same diagonal entries as $\mathbf{M}$. Let $\mathbf{I}_{d}$ represent the $d$ by $d$ identity matrix. For any two numbers $a,b\in\mathbb{R}$, we denote $a\wedge b:=\min\{a,b\}$ and $a\vee b:=\max\{a,b\}$. For any two sequences of positive numbers $\left\{a_{n}\right\}$ and $\left\{b_{n}\right\}$, we write $a_{n}\asymp b_{n}$ if $a_{n}=O\left(b_{n}\right)$ and $b_{n}=O\left(a_{n}\right)$. We write $b_{n}=\Omega\left(a_{n}\right)$ if $a_{n}=O\left(b_{n}\right)$, and $b_{n}=\Omega^{o}\left(a_{n}\right)$ if $b_{n}=\Omega\left(a_{n}\right)$ and $b_{n}\not\asymp a_{n}$. For any random variable $X\in\mathbb{R}$, we define the sub-Gaussian $\left(\|\cdot\|_{\psi_{2}}\right.$ ) and sub-exponential norms $\left(\|\cdot\|_{\psi_{1}}\right.$ ) of $X$ as follows:
$\|X\|_{\psi_{2}}:=\sup_{k\geq 1}k^{-1/2}\left(E|X|^{k}\right)^{1/k}\quad\text{ and }$
$\|X\|_{\psi_{1}}:=\sup_{k\geq 1}k^{-1}\left(E|X|^{k}\right)^{1/k}.$
Any $d$-dimensional random vector $X\in\mathbb{R}^{d}$ is said to be sub-Gaussian distributed with the sub-Gaussian constant $\sigma$ if
$\left\|\boldsymbol{v}^{T}\boldsymbol{X}\right\|_{\psi_{2}}\leq\sigma,\quad\text{ for any }\boldsymbol{v}\in\mathbb{S}^{d-1}.$
For any two vectors $\boldsymbol{v}_{1},\boldsymbol{v}_{2}\in\mathbb{R}^{d}$, let $\sin\angle\left(\boldsymbol{v}_{1},\boldsymbol{v}_{2}\right)$ be the sine of the angle between $\boldsymbol{v}_{1}$ and $\boldsymbol{v}_{2}$, with
$\left|\sin\angle\left(\boldsymbol{v}_{1},\boldsymbol{v}_{2}\right)\right|:=\sqrt{1-\left(\boldsymbol{v}_{1}^{T}\boldsymbol{v}_{2}\right)^{2}}.$
@@ 2 SPCA: Nonsparse Setting
where $U(\boldsymbol{x})=\boldsymbol{x}/||\boldsymbol{x}||I(\boldsymbol{x}\not=\boldsymbol{0})$ is the spatial sign function. By Theorem 4.4 in Oja (2010), we know that the eigenspace of the spatial sign covariance matrix ${\bf S}$ is identical to the eigenspace of the covariance matrix ${\bf\Sigma}$. Note that, according to Lemma B.1 in Han and Liu (2018), we know that the multivariate Kendall’s tau matrix. i.e.
@@
$\hat{\boldsymbol{\mu}}=\underset{\boldsymbol{\mu}\in\mathbb{R}^{p}}{\arg\min}\sum_{i=1}^{n}||\boldsymbol{X}_{i}-\boldsymbol{\mu}||_{2}.$ (4)
@@
Visuri et al. (2000) show that the influence functions of the sample spatial sign covariance matrix are uniformly bounded, indicating their robustness. In fact, the influence function of the sample spatial sign covariance matrix at a distribution $F$ symmetric around zero has the simple form
$IF(\boldsymbol{x},\hat{\bf S},F)=U(\boldsymbol{x})U(\boldsymbol{x})^{T}-\hat{{\bf S}}$
and is seen to be constant in the radius $\|\mathbf{x}\|$ of the contamination point $\boldsymbol{x}$.
@@
When the dimension $d$ is fixed and $rank({\bf\Sigma})=q$, Croux et al. (2002) showed the asymptotic normality of $\hat{\boldsymbol{u}}_{1}(\hat{\bf S})$, i.e.
$\sqrt{n}(\hat{\boldsymbol{u}}_{1}(\hat{\bf S})-\boldsymbol{u}_{1})\mathop{\rightarrow}\limits^{d}N(0,\sigma_{u1}^{2})$
where
$\sigma_{u1}^{2}=\sum_{l\neq 1}\left[\frac{\lambda_{l}({\bf\Sigma})\lambda_{1}({\bf\Sigma})b_{1l}}{\left(c_{1}\lambda_{1}({\bf\Sigma})-c_{l}\lambda_{l}({\bf\Sigma})\right)^{2}}\boldsymbol{u}_{l}\boldsymbol{u}_{l}^{T}\right]$
where
$c_{l}=E\left[u_{1}^{2}/\left(\gamma_{1}u_{1}^{2}+\ldots+\gamma_{q}u_{q}^{2}\right)\right]$
for $l=1,\ldots,q$, and for $1\leq j,l\leq q$
$b_{1l}=E\left[u_{1}^{2}u_{l}^{2}/\left(\gamma_{1}u_{1}^{2}+\ldots+\gamma_{q}u_{q}^{2}\right)\right]^{2}$
with $\left(u_{1},\ldots,u_{q}\right)$ the components of a random variable $\boldsymbol{u}$, uniformly distributed on the periphery of a unit sphere, and $\gamma_{1},\ldots,\gamma_{p}$ the standardized eigenvalues, that is $\gamma_{j}({\bf\Sigma})=\lambda_{j}({\bf\Sigma})/\left(\lambda_{1}({\bf\Sigma})+\ldots,\lambda_{q}({\bf\Sigma})\right)$.
To the best of our knowledge, as the dimension $d$ approaches infinity, there are currently no asymptotic results available for $\hat{\boldsymbol{u}}_{1}(\hat{\mathbf{S}})$. So we first analysis the convergence rate of $\boldsymbol{u}_{1}(\hat{\bf S})$ in high dimensional settings. According to Davis-Kahan inequality (Davis and Kahan, 1970; Wedin, 1972), to evaluate the convergence rate of $\boldsymbol{u}_{1}(\hat{\bf S})$ to $\boldsymbol{u}_{1}({\bf S})$, we first study the convergence rate of $\hat{{\bf S}}$ to ${\bf S}$ under the spectral norm.
Define $\zeta_{k}=E(r_{i}^{-k})$, $r_{i}=\|\boldsymbol{X}_{i}-\boldsymbol{\mu}\|$, $\nu_{i}=\zeta_{1}^{-1}r_{i}^{-1}$. We assume that
@@
The initial term in (6) bears a resemblance to the nonasymptotic bound for $\|\hat{{\bf K}}-{\bf K}\|$ presented in Theorem 3.1 of Han and Liu (2018). Here, $\hat{{\bf K}}$ represents the sample multivariate Kendall’s tau estimator. Essentially, this term constitutes the nonasymptotic bound for $\|\tilde{{\bf S}}-{\bf S}\|$, where $\tilde{{\bf S}}=\frac{1}{n}\sum_{i=1}^{n}U(\boldsymbol{X}_{i}-\boldsymbol{\mu})U(\boldsymbol{X}_{i}-\boldsymbol{\mu})^{T}$. This result is established using the matrix Bernstein inequality introduced by Tropp (2012). The second term in (6) arises from the convergence rate of the spatial median. A key distinction between the two estimators of ${\bf S}={\bf K}$, namely $\hat{{\bf K}}$ and $\hat{{\bf S}}$, lies in the necessity to estimate the location parameter for $\hat{{\bf S}}$. This additional step introduces complexity to the proof of its convergence rate. Conversely, the theoretical analysis of $\hat{{\bf K}}$ is simplified by obviating the need to estimate the location parameter. However, this simplification comes at the cost of increased computational burden, as $\hat{{\bf K}}$ is a second-order U-statistic, whereas $\hat{{\bf S}}$ is only a first-order U-statistic. Ultimately, if $\frac{r^{*}({\bf S})}{n^{\delta}\log d}\to 0$, the second term will be of smaller order compared to the first term.
@@
where $r^{**}(\boldsymbol{\Sigma}):=\|\boldsymbol{\Sigma}\|_{F}/\lambda_{1}(\boldsymbol{\Sigma})\leq\sqrt{d}$ is the “second-order” effective rank of the matrix $\boldsymbol{\Sigma}$. Additionally, if $\|{\bf\Sigma}\|_{F}\log d={\rm tr}({\bf\Sigma})$, we have $\lambda_{j}({\bf S})\asymp\lambda_{j}({\bf\Sigma})/{\rm tr}({\bf\Sigma})$. So
$\frac{\lambda_{1}(\mathbf{S})}{\lambda_{1}(\mathbf{S})-\lambda_{2}(\mathbf{S})}\asymp\frac{\lambda_{1}(\boldsymbol{\Sigma})}{\lambda_{1}(\boldsymbol{\Sigma})-\lambda_{2}(\boldsymbol{\Sigma})}.$
So, we can directly bound $\|\hat{\bf S}-{\bf S}\|_{2}$ and $\left|\sin\angle\left(\boldsymbol{u}_{1}(\widehat{\mathbf{S}}),\boldsymbol{u}_{1}(\mathbf{S})\right)\right|$ using ${\bf\Sigma}$.

### v2

@@ 1 Introduction
PCA (Principal Component Analysis) is a widely used statistical method for data dimensionality reduction. It transforms high-dimensional data into a lower-dimensional space while preserving as much of the original data’s variability as possible. PCA achieves this by identifying the directions of maximum variance in the data, known as principal components, and projecting the data onto these directions. This process removes redundant information and noise, making the data easier to handle and visualize. PCA is commonly applied in fields such as machine learning (Balcan et al., 2016), image processing (Chan et al., 2015), and finance (Lan and Du, 2019; Yang and Du, 2025), where dealing with high-dimensional data is common.
@@
Notation: Here we use the same notations as Han and Liu (2018). Let $\mathbf{M}=\left[\mathbf{M}_{jk}\right]\in\mathbb{R}^{d\times d}$ be a symmetric matrix and $\boldsymbol{v}=\left(v_{1},\ldots,v_{d}\right)^{T}\in\mathbb{R}^{d}$ be a vector. We denote $\boldsymbol{v}_{I}$ to be the subvector of $\boldsymbol{v}$ whose entries are indexed by a set $I$, and $\mathbf{M}_{I,J}$ to be the submatrix of $\mathbf{M}$ whose rows are indexed by $I$ and columns are indexed by $J$. We denote $\operatorname{supp}(v):=\left\{j:v_{j}\neq 0\right\}$. For $0<q<\infty$, we define the $\ell_{q}$ and $\ell_{\infty}$ vector norms as $\|\boldsymbol{v}\|_{q}:=\left(\sum_{i=1}^{d}\left|v_{i}\right|^{q}\right)^{1/q}$ and $\|\boldsymbol{v}\|_{\infty}:=\max_{1\leq i\leq d}\left|v_{i}\right|$. We denote $\|\boldsymbol{v}\|_{0}:=\operatorname{card}(\operatorname{supp}(\boldsymbol{v}))$. We define the matrix entry-wise maximum value and Frobenius norms as $\|\mathbf{M}\|_{\text{max }}:=\max\left\{\left|\mathbf{M}_{ij}\right|\right\}$ and $\|\mathbf{M}\|_{\mathrm{F}}=\left(\sum\mathbf{M}_{jk}^{2}\right)^{1/2}$. Let $\lambda_{j}(\mathbf{M})$ be the $j$ th largest eigenvalue of $\mathbf{M}$. If there are ties, $\lambda_{j}(\mathbf{M})$ is any one of the eigenvalues such that any eigenvalue larger than it has rank smaller than $j$, and any eigenvalue smaller than it has rank larger than $j$. Let $\boldsymbol{u}_{j}(\mathbf{M})$ be any unit vector $\boldsymbol{v}$ such that $\boldsymbol{v}^{T}\mathbf{M}\boldsymbol{v}=\lambda_{j}(\mathbf{M})$. Without loss of generality, we assume that the first nonzero entry of $\boldsymbol{u}_{j}(\mathbf{M})$ is positive. We denote $\|\mathbf{M}\|_{2}$ to be the spectral norm of $\mathbf{M}$ and $\mathbb{S}^{d-1}:=\left\{\boldsymbol{v}\in\mathbb{R}^{d}:\|\boldsymbol{v}\|_{2}=1\right\}$ to be the $d$-dimensional unit sphere. We define the restricted spectral norm $\|\mathbf{M}\|_{2,s}:=\sup_{\boldsymbol{v}\in\mathbb{S}^{d-1},\|\boldsymbol{v}\|_{0}\leq s}\left|\boldsymbol{v}^{T}\mathbf{M}\boldsymbol{v}\right|$, so for $s=d$, we have $\|\mathbf{M}\|_{2,s}=\|\mathbf{M}\|_{2}$. We denote $f(\mathbf{M})$ to be the matrix with entries $[f(\mathbf{M})]_{jk}=f\left(\mathbf{M}_{jk}\right)$. We denote $\operatorname{diag}(\mathbf{M})$ to be the diagonal matrix with the same diagonal entries as $\mathbf{M}$. Let $\mathbf{I}_{d}$ represent the $d$ by $d$ identity matrix. For any two numbers $a,b\in\mathbb{R}$, we denote $a\wedge b:=\min\{a,b\}$ and $a\vee b:=\max\{a,b\}$. For any two sequences of positive numbers $\left\{a_{n}\right\}$ and $\left\{b_{n}\right\}$, we write $a_{n}\asymp b_{n}$ if $a_{n}=O\left(b_{n}\right)$ and $b_{n}=O\left(a_{n}\right)$. We write $b_{n}=\Omega\left(a_{n}\right)$ if $a_{n}=O\left(b_{n}\right)$, and $b_{n}=\Omega^{o}\left(a_{n}\right)$ if $b_{n}=\Omega\left(a_{n}\right)$ and $b_{n}\not\asymp a_{n}$. For any random variable $X\in\mathbb{R}$, we define the sub-Gaussian $\left(\|\cdot\|_{\psi_{2}}\right.$ ) and sub-exponential norms $\left(\|\cdot\|_{\psi_{1}}\right.$ ) of $X$ as follows: $\|X\|_{\psi_{2}}:=\sup_{k\geq 1}k^{-1/2}\left(E|X|^{k}\right)^{1/k}\text{ and }\|X\|_{\psi_{1}}:=\sup_{k\geq 1}k^{-1}\left(E|X|^{k}\right)^{1/k}.$ Any $d$-dimensional random vector $X\in\mathbb{R}^{d}$ is said to be sub-Gaussian distributed with the sub-Gaussian constant $\sigma$ if $\left\|\boldsymbol{v}^{T}\boldsymbol{X}\right\|_{\psi_{2}}\leq\sigma,\text{ for any }\boldsymbol{v}\in\mathbb{S}^{d-1}.$ For any two vectors $\boldsymbol{v}_{1},\boldsymbol{v}_{2}\in\mathbb{S}^{d-1}$, let $\sin\angle\left(\boldsymbol{v}_{1},\boldsymbol{v}_{2}\right)$ be the sine of the angle between $\boldsymbol{v}_{1}$ and $\boldsymbol{v}_{2}$, with $\left|\sin\angle\left(\boldsymbol{v}_{1},\boldsymbol{v}_{2}\right)\right|:=\sqrt{1-\left(\boldsymbol{v}_{1}^{T}\boldsymbol{v}_{2}\right)^{2}}.$
@@ 2 SPCA: Nonsparse Setting
where $U(\boldsymbol{x})=\boldsymbol{x}/\|\boldsymbol{x}\|{{}_{2}}I(\boldsymbol{x}\not=\boldsymbol{0})$ is the spatial sign function. By Theorem 4.4 in Oja (2010), we know that the eigenspace of the spatial sign covariance matrix ${\bf S}$ is identical to the eigenspace of the covariance matrix ${\bf\Sigma}$. Note that, according to Lemma B.1 in Han and Liu (2018), we know that the multivariate Kendall’s tau matrix. i.e.
@@
$\hat{\boldsymbol{\mu}}=\underset{\boldsymbol{\mu}\in\mathbb{R}^{p}}{\arg\min}\sum_{i=1}^{n}\|\boldsymbol{X}_{i}-\boldsymbol{\mu}\|_{2}.$ (4)
@@
Visuri et al. (2000) show that the influence functions of the sample spatial sign covariance matrix are uniformly bounded, indicating their robustness. In fact, the influence function of the sample spatial sign covariance matrix at a distribution $F$ symmetric around zero has the simple form $IF(\boldsymbol{x},\hat{\bf S},F)=U(\boldsymbol{x})U(\boldsymbol{x})^{T}-\hat{{\bf S}}$ and is seen to be constant in the radius $\|\mathbf{x}\|{{}_{2}}$ of the contamination point $\boldsymbol{x}$.
@@
When the dimension $d$ is fixed and $rank({\bf\Sigma})=q$, Croux et al. (2002) showed the asymptotic normality of ${\boldsymbol{u}}_{1}(\hat{\bf S})$, i.e. $\sqrt{n}({\boldsymbol{u}}_{1}(\hat{\bf S})-{\boldsymbol{u}_{1}({\bf S})})\mathop{\rightarrow}\limits^{d}N(0,\sigma_{u1}^{2})$ where $\sigma_{u1}^{2}=\sum_{l\neq 1}\left[\frac{\lambda_{l}({\bf\Sigma})\lambda_{1}({\bf\Sigma})b_{1l}}{\left(c_{1}\lambda_{1}({\bf\Sigma})-c_{l}\lambda_{l}({\bf\Sigma})\right)^{2}}\boldsymbol{u}_{l}\boldsymbol{u}_{l}^{T}\right]$ where $c_{l}=E\left[u_{l}^{2}/\left(\gamma_{1}u_{1}^{2}+\ldots+\gamma_{q}u_{q}^{2}\right)\right]$ for $l=1,\ldots,q$, and for $1\leq j,l\leq q$, $b_{1l}=E\left[u_{1}^{2}u_{l}^{2}/\left(\gamma_{1}u_{1}^{2}+\ldots+\gamma_{q}u_{q}^{2}\right)\right]^{2}$ with $\left(u_{1},\ldots,u_{q}\right)$ the components of a random variable $\boldsymbol{u}$, uniformly distributed on the periphery of a unit sphere, and $\gamma_{1},\ldots,\gamma_{p}$ the standardized eigenvalues, that is $\gamma_{j}({\bf\Sigma})=\lambda_{j}({\bf\Sigma})/\left(\lambda_{1}({\bf\Sigma})+\ldots,\lambda_{q}({\bf\Sigma})\right)$.
To the best of our knowledge, as the dimension $d$ approaches infinity, there are currently no asymptotic results available for $\hat{\boldsymbol{u}}_{1}(\hat{\mathbf{S}})$. To fill this gap, we first analysis the convergence rate of $\boldsymbol{u}_{1}(\hat{\bf S})$ in high dimensional settings. According to Davis-Kahan inequality (Davis and Kahan, 1970; Wedin, 1972), to evaluate the convergence rate of $\boldsymbol{u}_{1}(\hat{\bf S})$ to $\boldsymbol{u}_{1}({\bf S})$, we first study the convergence rate of $\hat{{\bf S}}$ to ${\bf S}$ under the spectral norm.
Define $\zeta_{k}=E(r_{i}^{-k})$, $r_{i}=\|\boldsymbol{X}_{i}-\boldsymbol{\mu}\|{{}_{2}}$, $\nu_{i}=\zeta_{1}^{-1}r_{i}^{-1}$. We assume that
@@
The initial term in (6) bears a resemblance to the nonasymptotic bound for $\|\hat{{\bf K}}-{\bf K}\|{{}_{2}}$ presented in Theorem 3.1 of Han and Liu (2018). Here, $\hat{{\bf K}}$ represents the sample multivariate Kendall’s tau estimator. Essentially, this term constitutes the nonasymptotic bound for $\|\tilde{{\bf S}}-{\bf S}\|{{}_{2}}$, where $\tilde{{\bf S}}=\frac{1}{n}\sum_{i=1}^{n}U(\boldsymbol{X}_{i}-\boldsymbol{\mu})U(\boldsymbol{X}_{i}-\boldsymbol{\mu})^{T}$. This result is established using the matrix Bernstein inequality introduced by Tropp (2012). The second term in (6) arises from the convergence rate of the spatial median. A key distinction between the two estimators of ${\bf S}={\bf K}$, namely $\hat{{\bf K}}$ and $\hat{{\bf S}}$, lies in the necessity to estimate the location parameter for $\hat{{\bf S}}$. This additional step introduces complexity to the proof of its convergence rate. Conversely, the theoretical analysis of $\hat{{\bf K}}$ is simplified by obviating the need to estimate the location parameter. However, this simplification comes at the cost of increased computational burden, as $\hat{{\bf K}}$ is a second-order U-statistic, whereas $\hat{{\bf S}}$ is only a first-order U-statistic. Ultimately, if $\frac{r^{*}({\bf S})}{n^{\delta}\log d}\to 0$, the second term will be of smaller order compared to the first term.
@@
where $r^{**}(\boldsymbol{\Sigma}):=\|\boldsymbol{\Sigma}\|_{F}/\lambda_{1}(\boldsymbol{\Sigma})\leq\sqrt{d}$ is the “second-order” effective rank of the matrix $\boldsymbol{\Sigma}$. Additionally, if $\|{\bf\Sigma}\|_{F}\log d=o(1){\rm tr}({\bf\Sigma})$, we have $\lambda_{j}({\bf S})\asymp\lambda_{j}({\bf\Sigma})/{\rm tr}({\bf\Sigma})$ when $d\rightarrow\infty$. So,
$\frac{\lambda_{1}(\mathbf{S})}{\lambda_{1}(\mathbf{S})-\lambda_{2}(\mathbf{S})}\asymp\frac{\lambda_{1}(\boldsymbol{\Sigma})}{\lambda_{1}(\boldsymbol{\Sigma})-\lambda_{2}(\boldsymbol{\Sigma})},$
as $d\rightarrow\infty.$ So, we can directly bound $\|\hat{\bf S}-{\bf S}\|_{2}$ and $\left|\sin\angle\left(\boldsymbol{u}_{1}(\widehat{\mathbf{S}}),\boldsymbol{u}_{1}(\mathbf{S})\right)\right|$ using ${\bf\Sigma}$.